# Towards Understanding the Robustness of Diffusion-Based Purification: A Stochastic Perspective

### v2

@@ Appendix I Adopting VPSDE(DDPM++) and EDM Models in DBP
Our work holds significant potential for positive societal impacts across various sectors, including autonomous driving, facial recognition payment systems, and medical assistance. We are dedicated to enhancing the safety and trustworthiness of global AI applications. However, there are potential negative societal impacts, particularly concerning privacy protection, due to adversarial perturbations. Nonetheless, we believe that the positive impacts generally outweigh the potential negatives. Regarding the limitations, our approach could benefit from integrating insights from traditional adversarial training methods (Zhang et al., 2019; Shafahi et al., 2019; Wang et al., 2023), such as through more extensive data augmentation and a refined ADDT loss design. Nevertheless, these limitations are minor and do not significantly detract from the overall contributions of this paper. We believe that these new findings and perspectives could have a sustained impact on future research on DBP, which is a promising approach to adversarial defense and could be more valuable for real-world applications, although existing studies on DBP are at an early stage.

### v3

@@ Appendix I Adopting VPSDE(DDPM++) and EDM Models in DBP
Our work holds significant potential for positive societal impact across a range of sectors, including autonomous driving, facial recognition payment systems, and medical assistance. We are dedicated to enhancing the safety and trustworthiness of global AI applications. However, we acknowledge the potential negative societal impacts, particularly concerning privacy protection, due to adversarial perturbations. Despite this, we believe that the overall positive impacts outweigh these risks.
In terms of limitations, our approach could benefit from incorporating insights from traditional adversarial training methods (Zhang et al., 2019; Shafahi et al., 2019; Wang et al., 2023), such as through more extensive data augmentation and a refined ADDT loss design. Nevertheless, these limitations are relatively minor and do not diminish the overall contributions of this paper. We believe these new findings and perspectives will have a sustained impact on future research in DBP, which is a promising approach to adversarial defense and could prove even more valuable for real-world applications, despite the early stage of existing studies on DBP.